Please answer the following question in c++ language

IS INTENDED TO BE DONE BY HAND WITH POSSIBLE USE OF A CALCULATOR.   YOU NEED TO SHOW SOME WORK BECAUSE THE OBECTIVE IS KNOWING WHAT A PROGRAM IS DOING, NOT RUNNING A PROGRAM. SHOWING HOW YOU GOT ANSWERS IS REQUIRED.

The divide and conquer Fast Fourier Transform (FFT) has run time of order   

a) O(logn)                      

b) O(n logn)  

c) O( )  

d) O(n)
Answer #1

Answer is

B)O(nlogn)

It is fast, and has running time in Θ(n log n) • There are many variations, often to take advantage of the numerical structure of n. Most commonly n is a power of 2. It can also be computed optically using lenses.

A Recursive Fast Fourier Transform #basis case Initialize results to 0. • Divide - even subproblem . - odd subproblem # FOCUSAn N log N algorithm Like in merge sort, in each recursive call, we divide the number of elements in half. The number of leve

Recursive FFT FFT(n,(a, a, .... a -1]): if n=1: return ao Feven = FFT(n/2, (a, a, ..., an-2]) Fodd = FFT(n/2, [a, a3, ..., an